OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44

OpenStudy (anonymous):

On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (54 + x)32 + (82 - (54 + x)) ∙ 44 = C (54 - x)32 + (82 - 54) ∙ 44 = C (82 + x)32 + (82 - (54 + x)) ∙ 44 = C (82 - x)32 + (82 - 54) ∙ 44 = C